OpenStudy (anonymous):

a body of mass 2 kg is suspended in an elevator by means of a spring balance the balance reads its weight when elevator moves up with an acceleration of 5ms_2 as 30.5n b 29.6n c 26.5n d 9.8n

OpenStudy (anonymous):

this question is a bit lengthy....

OpenStudy (anonymous):

the actual weight of the mass is 20 newtons......first of all..

OpenStudy (anonymous):

when the elevator moves upward with an acceleration a = 5 m/s^2 ......then the spring balance will show(give reading) of apparent weight more than the actual weight ..... which means that the choice "d"----9.8 is definitely wrong....

OpenStudy (anonymous):

now while the elevator is moving upward with an acceleration ...... there are two forces acting on it ......the upward tension "T" and downward weight "w".....

OpenStudy (anonymous):

As the mass "m" along with the elevator are accelerating upward with acceleration "a".....So a net force F=ma is acting in the upward direction ..... this net force F=ma= the difference between T and w. T>W because the elevator+mass are both accelerating upward.

OpenStudy (anonymous):

Therefore, F=ma=T-W || ma=T-W Putting values|| ma=T-mg|| T=ma+mg|| T=m(a+g)|| T=2(5+9.8)|| T=29.6N As the tension is equal to the apparent weight of the mass being shown by the spring balance....

OpenStudy (anonymous):

So option b is the right choice....

OpenStudy (anonymous):

Did you understand?